OpenStudy (anonymous):

What is the value of the expression? –15.6 + (–3.4 + 8) A. –27 B. –26 C. –11 D. –10

OpenStudy (anonymous):

i say its a?

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Solve from the parenthesis.

OpenStudy (anonymous):

k hold on

OpenStudy (anonymous):

–3.4 + 8= ?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

4.2 @Kamizamurai

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

–15.6 + (–3.4 + 8) =? –15.6+(4.6) =?

OpenStudy (anonymous):

= -11

OpenStudy (anonymous):

oh srry i was away 4 a while and thx

OpenStudy (anonymous):

Np :)